OpenStudy (anonymous):

trail and error?!? n^2 - 16n + 48

OpenStudy (anonymous):

=0?

OpenStudy (amistre64):

make a guess :) then adjust it

OpenStudy (anonymous):

i got (n - 4) (n - 12)

OpenStudy (amistre64):

if one guess produces a postive value, and another produces a negative value; then somewhere between those guesses is a root

OpenStudy (anonymous):

I prefer to do it like that: \[n^2-16n+48=0\]\[(x-f)(x-g)=(x^2+fg-fx-gx)\]Therefore \[fg=48\]\[f+g=-16\]Now its easier to guess. There is also a formula to solve directly, but since you need to guess...

OpenStudy (anonymous):

Oh, hahaha didnt see you already had the answer

OpenStudy (anonymous):

And its +16 on my answer not -16